Verify that the vector Xa is a particular solution of the given nonhomogeneous linear system. ,_21 et+ 1 tef, one has 1 te isdx dt x(t), y(t)) Need Help?Read It Watch It Talk to a Tutor
Answer #1

Given system of diff eq is

x = x + 2y

To solve above diff eq , first will find eigen values of coef matrix and corresponding eigen vector .

det C 2 (1-c)(3-c)--8=0 (c + 1)(c-5) = 0

Let u be an eigen vector corr to -1,

3 4

:1 310 Lu 4 3+1

2 12 24

Let v be an eigen vector corr to eigen value 5,

ひひ 2 3 4

\begin{bmatrix} 1-5& 2\\ 4 & 3-5 \end{bmatrix}\begin{bmatrix} v_1\\ v_2\end{bmatrix}=0

2 ひひ 2 4

201 V)

Hence general solution is

2

                 AeBe Ae-2Bes

x(t)= -Ae^{-t}+Be^{5t},y(t)=Ae^{-t}-2Be^{5t}